the weight of a dart to be used in a tournament is normally distributed with a mean of 17gms and a standard deviation of 0.3gms. find the probability that a dart choose at random has it weight between 16.55gms and 17.45gms 


1
Expert's answer
2021-12-07T12:38:18-0500

"z=\frac{x-\mu}{\sigma}"


"z_1=\frac{16.55-17}{0.3}=-1.5"


"z_2=\frac{17.45-17}{0.3}=1.5"


"P(16.55<x<17.45)=P(z_1<z<z_2)=P(z<1.5)-P(z<-1.5)="


"=0.9332-0.0668=0.8664"